OpenStudy (anonymous):

sqrt (16x- 16) + sqrt (25x^3- 25x^2)

OpenStudy (anonymous):

my teacher told us not to like mess with the binomials so im confused

OpenStudy (anonymous):

You can factor coefficients out of the first inside expression, so you get: sqrt(16(x-1)) - sqrt(25x^2(x-1)). The 16 in the first term comes to be a 4 on the outside. The 25x^2 inside the second term comes to be a 5x on the outside. 4sqrt(x-1)-5x*sqrt(x-1) And you can simplify by factoring out the sqrt(x-1): sqrt(x-1)*(4-5x).

OpenStudy (anonymous):

im so confused

OpenStudy (anonymous):

Do you get that (16x-16)=16(x-1)?

OpenStudy (anonymous):

o k now i get it wen u put it simplerr

OpenStudy (anonymous):

But 16 is a square so you can bring out it's root

OpenStudy (anonymous):

but its connected to a binomial so wich one of them...or both and 25 is also a square connected to a binomial in the laast half of the problem

OpenStudy (anonymous):

Well, we rewrite \(\sqrt{16(x-1)}\) as \(\sqrt{16}\sqrt{x-1}\) and simplify.

OpenStudy (anonymous):

Can you rewrite the second half now?

OpenStudy (anonymous):

yse so the answer is 5x + 4 sqrt x-1?

OpenStudy (anonymous):

(4-5x)

OpenStudy (anonymous):

and i knock off the sqrt x-1?

OpenStudy (anonymous):

no, no, no, just replace the 5x+4 with (4-5x)

OpenStudy (anonymous):

o okay

OpenStudy (anonymous):

do u mean 4 + 5x